OpenStudy (anonymous):

3w-10w = ?? Please give how you did the question too-not just the answer! Thanks so much!

OpenStudy (anonymous):

oh i have answer choices too--- 13w 7w -7w -7

OpenStudy (anonymous):

-7w

OpenStudy (anonymous):

3w-10w=w(3-10)=-7w

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

Oh lol thanks! :D

OpenStudy (anonymous):

lol kid